express as the cofunction of a positive acute angle not using a calculator of sin 340 degress.
Answer by venugopalramana(3286)

SIN(340)=SIN(360-20)=-SIN(20)...SINCE SIN IS NEGATIVE IN IV QUADRANT..SIN(360-X)=-SIN(X)